### v1

# Mass gap in non-perturbative quadratic $\mathcal{R}^{2}$ gravity via Dyson-Schwinger

Abstract
We apply in a simple model derived from quadratic $\mathcal{R}^{2}$ gravity the technique of Dyson-Schwinger equations to solve for its corresponding quantum theory. Particularly, we solve the classical equations of motion to get a solution to the hierarchy of Dyson-Schwinger equations in the limit of large coupling. The model is obtained by looking for conformally flat solutions. The model admits Higgs-like solutions with a single particle having a finite mass. Assuming a Fubini-Lipatov instanton for the vacuum instead, the spectrum is given by an infinite set of excitations of finite mass. In both cases, the model develops a mass gap due to the presence of the quadratic Ricci scalar, after the breaking of the conformal symmetry.
I Introduction
One of the most successful theories of the past century and certainly the most successful theory in the context of gravity is Albert Einstein’s general relativity (GR). This theory has stood the test of time again and again which has been verified by several observations. These experimental observations range from those as recent as the discovery of gravitational waves, with coalescing binary black hole astrophysical systems LIGOScientific:2016aoc , to the extraordinary optical images of the black holes in the center of our galaxy and the M87 one as captured by the Event Horizon Telescope EventHorizonTelescope:2019dse ; EventHorizonTelescope:2019ggy ; EventHorizonTelescope:2022xnr ; EventHorizonTelescope:2022wok ; EventHorizonTelescope:2022xqj . Extending GR, the theory suitably can account for matter fields, and forms the edifice of our present understanding of modern day cosmology, including the early universe right from the origin of hot big bang.
Even with so much success at large scales, nonetheless, at the quantum frontier GR is known to be non-renormalizable by perturbative methods Goroff:1985sz ; Goroff:1985th and may not remain perturbative at length scales much below the Planck one. Also it encounters Big Bang and Black Hole singularity problems and known to pathological behaviour in the ultra-violate (UV) Salvio:2018crh . GR also suffers from the what is known as the “conformal-factor problem”: in simple terms this is nothing but absence of a consistent Euclidean path integral. Moreover this becomes quite serious since all known quantum field theories (QFT) can be consistently defined only as analytic continuations of the corresponding Euclidean version, see Ref. Salvio:2024joi for details.
The issue of non-renormalizability and the issue of conformal-factor is solved when GR is augmented with quadratic-in-curvature terms which is often called quadratic gravity (See Salvio:2018crh ; Donoghue:2021cza for reviews.). It has been shown to be renormalizable Weinberg:1974tw ; Deser:1975nv ; Stelle:1976gc ; Barvinsky:2017zlx and following the prescription of Ref. Gibbons:1994cg one may determine the Euclidean action which implies that it is also free from the conformal-factor problem in most of the regions of its parameter space Menotti:1989ms . Such quadratic gravity theories also address the singularity issues. The classical Hamiltonian of quadratic gravity is unbounded from below, something known in the literature as the Ostrogradsky’s instability issue as shown in Ref. ostro . Frankly, this is common in all kinds of higher-derivative gravity extensions. Nonetheless estimations of runaway rates show that such a decay can be slow enough to be compatible with observations as studied in Refs. Salvio:2019ewf ; Gross:2020tph ; Held:2021pht ; Held:2023aap .
In this paper we offer a non-perturbative formulation of quadratic $\mathcal{R}^{2}$ gravity. This is based on the Dyson-Schwinger approach and may certainly provide us with a fully quantum and non-perturbative justification for these classical cosmological findings in the long run. Furthermore, such a formulation that we show here can also be utilised to validate and check consistency with other independent perturbative methods and sometimes even background-dependent approaches to formulate quadratic gravity which been shown to preserve unitarity and (meta)stablity. In the context of renormalizable theories, these approaches include the Lee-Wick Lee:1969fy ; Salvio:2018kwh ; Donoghue:2019fcb ; Holdom:2021hlo the fakeon prescription Anselmi:2017ygm and string-inspired non-local ghost-free approaches to higher-derivative gravity Frasca:2020jbe ; Frasca:2020ojd ; Frasca:2021iip ; Frasca:2022duz ; Frasca:2022gdz . Particularly for quadratic gravity this has been studied rigorously in Ref. Salvio:2024joi .
The novel tool to treat the non-perturbative systems in the framework of Dyson-Schwinger Equations that we propose in this paper are based on the exact solution of the background equations of motion in terms of Jacobi elliptic functions. This is inspired by the analytic approach of Dyson-Schwinger equations (DSE), originally devised by Bender, Milton and Savage Bender:1999ek . Using this we are able to represent the Green’s functions of the theory purely analytically, therefore we understand the effect of the background on the interactions that remains to be valid even in the strong coupling regimes Frasca:2015yva . This approach has been widely used to investigate QCD with great success with respect to results obtain closely matching with lattice QCD results Frasca:2021yuu ; Frasca:2021mhi ; Frasca:2022lwp ; Frasca:2022pjf ; Chaichian:2018cyv and to the scalar sector Frasca:2015wva  [^1] Moreover extensions beyond the traditional QFT and SM of particle physics to other types of models involving new gauge sectors and string-inspired non-local theories have been widely discussed using this methodology Frasca:2019ysi ; Chaichian:2018cyv ; Frasca:2017slg ; Frasca:2016sky ; Frasca:2015yva ; Frasca:2015wva ; Frasca:2013tma ; Frasca:2012ne ; Frasca:2009bc ; Frasca:2010ce ; Frasca:2008tg ; Frasca:2009yp ; Frasca:2008zp ; Frasca:2007uz ; Frasca:2006yx ; Frasca:2005sx ; Frasca:2005mv ; Frasca:2005fs . In context to experiments concerning particle physics phenomenology like non-perturbative hadronic vacuum polarization contributions to the muon anomalous magnetic moment $(g-2)_{\mu}$ was predicted in Ref.Frasca:2021yuu , QCD in the non-perturbative regimes Frasca:2021mhi ; Frasca:2022lwp ; Frasca:2022pjf , Higgs-Yukawa theory Frasca:2023qii , finite temperature field theory Frasca:2023eoj , non-perturbative false vacuum transitions leading to phase transitions Frasca:2022kfy ; Calcagni:2022tls ; Calcagni:2022gac , dark energy Frasca:2022vvp , and investigations of the mass gap and confinement in string-inspired infinite-derivative and Lee-Wick theories Frasca:2020jbe ; Frasca:2020ojd ; Frasca:2021iip ; Frasca:2022duz ; Frasca:2022gdz . These tools were also employed to implement dynamical generation electroweak (EW) scale via scalar, vector boson and fermionic condensates too very recently Frasca:2024fuq ; Chatterjee:2024dgw ; Frasca:2024pmv ; Frasca:2023eoj ; Frasca:2024ame as an approach to address the gauge hierarchy problem in the Standard Model (SM) of particle physics.
A final formulation of the theory should involve quantum cosmology with a quantum mechanical “wave function of the universe”, as investigated by Hawking and Hartle Hartle:1983ai . This actually can be used to solve the Schrodinger equation, also known as Wheeler-DeWitt equation in this context DeWitt:1967yk ; Wheeler . Therefore a fully consistent quantum cosmology with GR extended to quadratic gravity is what we allude to, since this theory contains Starobinsky’s model of inflation (also known as $\mathcal{R}^{2}$ inflation) Starobinsky:1980te , which is currently in perfect agreement with the cosmic microwave background radiation (CMBR) observations assuming the cosmic inflationary paradigm Ade:2015lrj ; Planck2018:inflation ; BICEP:2021xfz .
Yet another application of a non-perturbative quadratic gravity via Dyson-Schwinger approach would be in the paradigm of asymptotic safety originally proposed by Steve Weinberg WeinbergAS , where gravity is made ultra-violate (UV) complete and shown to be renormalizable due to presence of a non-perturbative UV fixed point. Investigations of such fixed points have been extensively carried out for quadratic $\mathcal{R}^{2}$ gravity in Refs. Benedetti:2009rx ; Falls:2020qhj which lead to consistent realization of Asymptotic Safe Gravity.
As we will show that the Einstein-Hilbert term plus a quadratic term of the Ricci scalar, for strong fields. Conformal symmetry is broken by the appearance of a mass gap in the theory. The mass spectrum of theory is a discrete spectrum with an infinite tower of harmonic-oscillator like excitations. The mass gap scale is proportional to the Starobinsky mass which is the coefficient of the $\mathcal{R}^{2}$ term which represents the self-interactions among the gravitons and the scalar degrees of freedom.
The paper is organised as follows: In Sec. II, we introduce the model and derive the scalar field theory arising from it. In Sec. III, we discuss the classical solutions for the scalar field theory. In Sec. IV, we quantize the model ans solve the set of Dyson-Schwinger equations both with a null or non-null the cosmological constant. In Sec.V, we provide the conclusions.
II Scalar Model for Quadratic Starobinsky gravity: A small review
We present a scalar field model from the full action of the quadratic gravity Starobinsky:1980te ; Mukhanov:1981xt ; Kaneda:2010ut and then we will quantize it with the technique of the Dyson-Schwinger equations.
$S=-\frac{M_{p}^{2}}{2}\int d^{4}x\left(\mathcal{R}-2\Lambda+\frac{\mathcal{R}^{2}}{6M^{2}}\right),$ (1)
where $M_{p}$ is the Planck mass and $M$ is a new mass scale in this theory which is called the Starobinsky mass or also known as the conformal coupling. This action can be re-written in terms of the interaction between gravity and a scalar field $\chi$ after performing the transformation Whitt:1984pd ; Kehagias:2013mya
$g_{\mu\nu}\rightarrow e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}g_{\mu\nu}.$ (2)
which gives:
$S=\int{\rm d}^{4}x\sqrt{-g}\left[-\frac{M_{\rm p}^{2}}{2}(\mathcal{R}-2\Lambda)+\frac{1}{2}g^{\mu\nu}\partial_{\mu}\chi\partial_{\nu}\chi+\frac{3}{4}M_{\rm p}^{2}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}\right)^{2}\right].$ (3)
This action yields the following sets of equations of motion, which are given by $\frac{\delta S}{\delta g_{\mu\nu}}$ and $\frac{\delta S}{\delta\chi}$:
$\mathcal{R}_{\mu\nu}=-M_{p}^{-2}\left[\partial_{\mu}\chi\partial_{\nu}\chi-g_{\mu\nu}\left(\partial_{\sigma}\chi\partial^{\sigma}\chi+\frac{3}{2}M_{\rm p}^{2}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}\right)^{2}\right)-2g_{\mu\nu}M_{p}^{2}\Lambda\right]$
$\frac{1}{\sqrt{-g}}\partial_{\mu}g^{\mu\nu}\sqrt{-g}\partial_{\nu}\chi=-\sqrt{\frac{3}{2}}M_{\rm p}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}\right)e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}.$ (4)
Now after removing the term $\partial_{\mu}\chi\partial_{\nu}\chi$ by the proper gauge choice in harmonic coordinates (see Appendix A) and assuming a Ricci flat metric, as per Minkowski metric, we get the following set of simplified equations:
$\partial_{\sigma}\chi\partial^{\sigma}\chi+\frac{3}{2}M_{\rm p}^{2}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}\right)^{2}+2M_{p}^{2}\Lambda=0$
$\frac{1}{\sqrt{-g}}\partial_{\mu}g^{\mu\nu}\sqrt{-g}\partial_{\nu}\chi=-\sqrt{\frac{3}{2}}M_{\rm p}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}\right)e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}.$ (5)
This yields an overdetermined system of equations, assuming the manifold has a constant determinant as we are considering conformally flat metrics,
$\partial_{\sigma}\chi\partial^{\sigma}\chi+\frac{3}{2}M_{\rm p}^{2}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}\right)^{2}+2M_{p}^{2}\Lambda=0$
$\Box\chi=-\sqrt{\frac{3}{2}}M_{\rm p}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}\right)e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}.$ (6)
For $\chi\gg M_{p}$, the system is consistent yielding a solution:
$\chi_{\infty}=p_{\mu}x^{\mu}+\theta,$ (7)
where $p_{\mu}$ and $\theta$ are the appropriate integration constants, $x^{\mu}$ space-time coordinates. This yields $p^{2}+(3/2)M_{\rm p}^{2}M^{2}+2M_{\rm p}^{2}\Lambda=0$ for the determination of the constants $p_{\mu}$ Anyway, if we assume universe of be isotropic, this solution just recover the well-known (Anti)de Sitter solution in conformal coordinates if we assume $\Lambda<0$ and $|\Lambda|>M^{2}$, otherwise, we can neglect this solution as unphysical. Taking a dimensionless rescaled field variable $\varphi=\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}$, rescaled the dimensionful field $\chi$ with respect to the Planck mass we turn back to the full set of eq.(II), using which we further get:
$\partial_{\sigma}\varphi\partial^{\sigma}\varphi+M^{2}\left(1-e^{-\varphi}\right)^{2}+\frac{4}{3}\Lambda=0$
$\Box\varphi=-M^{2}\left(1-e^{-\varphi}\right)e^{-\varphi},$ (8)
with the Dálembertain operator, $\Box=\partial_{t}^{2}-\Delta_{2}$. Here in this computation $\partial_{t}$ and $\partial_{\sigma}$ denote the partial differentiation with respect to $t$ and $\sigma$ respectively. Let us take $\varphi=-\log Z$, we get
$Z^{-2}\partial_{\sigma}Z\partial^{\sigma}Z+M^{2}\left(1-Z\right)^{2}+\frac{4}{3}\Lambda=0$
$-Z^{-1}\Box Z+Z^{-2}\partial_{\sigma}Z\partial^{\sigma}Z=-M^{2}\left(1-Z\right)Z.$ (9)
The insert the first equation into to second obtaining
$-Z^{-1}\Box Z-M^{2}\left(1-Z\right)^{2}-\frac{4}{3}\Lambda=-M^{2}(1-Z)Z.$ (10)
$-Z^{-1}\Box Z-M^{2}\left(1-2Z+Z^{2}\right)-\frac{4}{3}\Lambda=-M^{2}(Z-Z^{2}).$ (11)
$-Z^{-1}\Box Z=-M^{2}(3Z-2Z^{2})+M^{2}+\frac{4}{3}\Lambda.$ (12)
$\Box Z=3M^{2}Z^{2}-2M^{2}Z^{3}-M^{2}Z-\frac{4}{3}\Lambda Z.$ (13)
Thus, we are left with a quartic scalar $\phi^{4}$-like theory which can be easily seen upon the introduction of the following mapping of variables:
$Z=\frac{1}{\sqrt{2}M}\phi+\frac{1}{2},$ (14)
that gives us
$\Box\phi=-\frac{2\sqrt{2}}{3}M\Lambda+\left(-\frac{4}{3}\Lambda+\frac{1}{2}M^{2}\right)\phi-\phi^{3}.$ (15)
Such a potential could signal a false vacuum decay yielding, after integrating on $\phi$ and changed sign to the rhs of eq.(15),
$V(\phi)=\frac{2\sqrt{2}}{3}M\Lambda\phi-\frac{1}{2}\left(-\frac{4}{3}\Lambda+\frac{1}{2}M^{2}\right)\phi^{2}+\frac{1}{4}\phi^{4}.$ (16)
We have set the integration constant to 0.
Figure 1: Plot of the potential of the conformal factor. It is seen the possibility of false vacuum decay for $3M^{2}/16>\Lambda$. All the physical quantity are dimensionless being scaled with the Planck mass $M_{P}$.
Fig. 1 shows different potentials varying the Starobinsky mass $M$ with respect to the Planck mass for a given value of the cosmological constant. The characteristic shape of a false vacuum decay potential can be recognized.
III Classical solutions
III.1 $\Lambda\neq 0$
We can solve eq.(15) when we consider the term $-\frac{2\sqrt{2}}{3}M\Lambda$ as a perturbation. Thus, we write
$\phi(x)=\sum_{n=0}^{\infty}\epsilon^{n}\phi_{n}(x)$ (17)
where we have introduced a fictitious parameter $\epsilon$ for bookkeeping the order. We will set $\epsilon=1$ at the end of computation. We will get
$\sum_{n=0}^{\infty}\epsilon^{n}\Box\phi_{n}(x)=-\epsilon\frac{2\sqrt{2}}{3}M\Lambda+\mu_{0}^{2}\sum_{n=0}^{\infty}\epsilon^{n}\phi_{n}(x)-[\phi_{0}^{3}+3\epsilon\phi_{1}\phi_{0}^{2}+\epsilon^{2}(3\phi_{1}^{2}\phi_{0}+3\phi_{0}^{2}\phi_{2})+\epsilon^{3}(\phi_{1}^{3}+6\phi_{0}\phi_{1}\phi_{2}+3\phi_{0}^{2}\phi_{3})+\ldots].$ (18)
Here $\mu_{0}^{2}=-\frac{4}{3}\Lambda+\frac{1}{2}M^{2}$. This will yield the set of equations
$\Box\phi_{0}$ $=$ $\mu_{0}^{2}\phi_{0}-\phi_{0}^{3},$
$\Box\phi_{1}$ $=$ $\mu_{0}^{2}\phi_{1}-3\phi_{0}^{2}\phi_{1}-\frac{2\sqrt{2}}{3}M\Lambda,$
$\Box\phi_{2}$ $=$ $\mu_{0}^{2}\phi_{2}-(3\phi_{1}^{2}\phi_{0}+3\phi_{0}^{2}\phi_{2}),$
$\Box\phi_{3}$ $=$ $\mu_{0}^{2}\phi_{3}-(\phi_{1}^{3}+6\phi_{0}\phi_{1}\phi_{2}+3\phi_{0}^{2}\phi_{3}),$ (19)
$\vdots$ .
We get the following solutions (this class of solutions arises from our preceding studies for different field theories Frasca:2015yva ; Frasca:2021yuu ; Frasca:2021mhi ; Frasca:2022lwp ; Chaichian:2018cyv ; Frasca:2022vvp ; Frasca:2015wva ; Frasca:2017slg )
$\phi_{0}(x)=\sqrt{\frac{2}{3}}|\mu_{0}|\text{dn}\left(p\cdot x+\theta,-1\right),$ (20)
where dn is a Jacobian elliptic function and having fixed the dispersion relation to
$p^{2}=\frac{1}{3}\mu_{0}^{2},$ (21)
and where $\theta$ is an arbitrary constant arising from integration. To solve for $\phi_{1}$, we need the Green function $C_{1}$ given by the following Lamé equation
$\Box C_{1}(x,x^{\prime})=(\mu_{0}^{2}-3\phi_{0}^{2}(x))C_{1}(x,x^{\prime})+\delta^{4}(x-x^{\prime}).$ (22)
We can get a straightforward solution by noting that the homogeneous equation has the solution
$C_{1h}(x)=\text{cn}\left(p\cdot x+\theta,-1\right)\text{sn}\left(p\cdot x+\theta,-1\right),$ (23)
where cn and sn are Jacobian elliptic functions. Thus, using the results given in Ref. Frasca:2023uaw , we obtain in momentum space
$C_{1}(p)=\frac{\sqrt{2}\pi^{3}}{K^{3}(-1)}\sum_{n=1}^{\infty}n^{2}\frac{e^{-n\pi}}{1+e^{-2n\pi}}\frac{1}{p^{2}+m_{n}^{2}},$ (24)
where $K(-1)$ is the complete elliptic integral of the first kind and
$m_{n}=n\frac{\pi}{K(-1)}\frac{|\mu_{0}|}{\sqrt{3}}.$ (25)
It should be noticed the presence of a zero mass excitations in the spectrum but it is not propagating as $C_{1}(p)$ ha a null contribution in this case. This applies to our discussion for the quantum case. Thus,
$\phi_{1}(x)$ $=$ $-\frac{2\sqrt{2}}{3}M\Lambda\int d^{4}x^{\prime}C_{1}(x-x^{\prime}),$
$\phi_{2}(x)$ $=$ $-3\left(\frac{2\sqrt{2}}{3}M\Lambda\right)^{2}\int d^{4}x^{\prime}C_{1}(x-x^{\prime})\left[\int d^{4}x^{\prime\prime}C_{1}(x^{\prime}-x^{\prime\prime})\right]^{2}\phi_{0}(x^{\prime}),$
$\phi_{3}(x)$ $=$ $-\left(\frac{2\sqrt{2}}{3}M\Lambda\right)^{3}\left[\int d^{4}x^{\prime}C_{1}(x-x^{\prime})\left[\int d^{4}x^{\prime\prime}C_{1}(x^{\prime}-x^{\prime\prime})\right]^{3}\right.$ (26)
$\left.+18\int d^{4}x^{\prime}C_{1}(x-x^{\prime})\phi_{0}(x^{\prime})\int d^{4}x^{\prime\prime}C_{1}(x^{\prime}-x^{\prime\prime})\int d^{4}x^{\prime\prime\prime}C_{1}(x^{\prime\prime}-x^{\prime\prime\prime})\left[\int d^{4}x^{\prime\prime\prime\prime}C_{1}(x^{\prime\prime\prime}-x^{\prime\prime\prime\prime})\right]^{2}\phi_{0}(x^{\prime\prime\prime\prime})\right],$
$\vdots$ .
III.2 $\Lambda=0$
For $\Lambda=0$, we need to introduce an arbitrary source $j$ and compute the functional series
$\phi(x)=\phi_{0}(x)+\sum_{n=1}^{\infty}\int C_{n}(x,x_{1},\ldots,x_{n-1})\prod_{k=1}^{n}j(x_{k})d^{4}x_{k},$ (27)
where we are assuming $\phi$ a functional of an arbitrary source $j$ and
$C_{n}(x,x_{1},\ldots,x_{n-1})=\left.\frac{\delta^{n}\phi(x)}{\delta j(x_{1})\ldots\delta j(x_{n-1})}\right|_{j=0}.$ (28)
By iterating the derivation with respect to $j$ and starting from the equation of motion
$\Box\phi(x)=\frac{M^{2}}{2}\phi(x)-\phi^{3}(x)+j(x),$ (29)
we get
$\Box\phi_{0}(x)$ $=$ $\frac{M^{2}}{2}\phi_{0}(x)-\phi_{0}^{3}(x),$
$\Box C_{1}(x,x_{1})$ $=$ $\left(\frac{M^{2}}{2}-3\phi_{0}^{2}(x)\right)C_{1}(x,x_{1})+\delta^{4}(x-x_{1}),$
$\Box C_{2}(x,x_{1},x_{2})$ $=$ $\left(\frac{M^{2}}{2}-3\phi_{0}^{2}(x)\right)C_{2}(x,x_{1},x_{2})-6\phi_{0}(x)C_{1}(x,x_{2})C_{1}(x,x_{1}),$
$\Box C_{3}(x,x_{1},x_{2},x_{3})$ $=$ $\left(\frac{M^{2}}{2}-3\phi_{0}^{2}(x)\right)C_{3}(x,x_{1},x_{2},x_{3})-6C_{1}(x,x_{1})C_{1}(x,x_{2})C_{1}(x,x_{3})-6\phi_{0}(x)C_{2}(x,x_{1},x_{2})C_{1}(x,x_{3})$ (30)
$-6\phi_{0}(x)C_{1}(x,x_{1})C_{2}(x,x_{2},x_{3})-6\phi_{0}(x)C_{2}(x,x_{1},x_{3})C_{1}(x,x_{2}),$
$\vdots$ .
We can use the results of the preceding section to solve this set of equations yielding
$\phi_{0}(x)=\frac{1}{\sqrt{3}}M\text{dn}\left(p\cdot x+\theta,-1\right),$ (31)
where one can see M as the Starobinsky mass in the original Lagrangian 1, dn is a Jacobian elliptic function and
$p^{2}=\frac{1}{6}M^{2}.$ (32)
Similarly, also in this case
$C_{1}(p)=\frac{\sqrt{2}\pi^{3}}{K^{3}(-1)}\sum_{n=1}^{\infty}n^{2}\frac{e^{-n\pi}}{1+e^{-2n\pi}}\frac{1}{p^{2}+m_{n}^{2}},$ (33)
where $K(-1)$ is the complete elliptic integral of the first kind and
$m_{n}=n\frac{\pi}{K(-1)}\frac{M}{\sqrt{6}}.$ (34)
We can iterate obtaining
$C_{2}(x,x_{1},x_{2})$ $=$ $-6\int d^{4}x^{\prime}C_{1}(x,x^{\prime})\phi_{0}(x^{\prime})C_{1}(x^{\prime},x_{2})C_{1}(x^{\prime},x_{1}),$
$C_{3}(x,x_{1},x_{2},x_{3})$ $=$ $-6\int d^{4}x^{\prime}C_{1}(x,x^{\prime})C_{1}(x^{\prime},x_{1})C_{1}(x^{\prime},x_{2})C_{1}(x^{\prime},x_{3})-6\int d^{4}x^{\prime}C_{1}(x,x^{\prime})\phi_{0}(x^{\prime})C_{2}(x^{\prime},x_{1},x_{2})C_{1}(x^{\prime},x_{3})$ (35)
$-6\int d^{4}x^{\prime}C_{1}(x,x^{\prime})\phi_{0}(x^{\prime})C_{1}(x^{\prime},x_{1})C_{2}(x^{\prime},x_{2},x_{3})-6\int d^{4}x^{\prime}C_{1}(x,x^{\prime})\phi_{0}(x^{\prime})C_{2}(x^{\prime},x_{1},x_{3})C_{1}(x^{\prime},x_{2}),$
$\vdots$ .
We see that our solution is completely obtained by knowing $\phi_{0}$ and the Green function $C_{1}(x,x^{\prime})$ that is translation invariant and can be written as $C_{1}(x-x^{\prime})$. We will see that the same result can be obtained in quantum field theory providing a Gaussian solution to the set of Dyson-Schwinger equations of the theory.
It is also possible to get a constant solution at leading order
$\phi_{0}=\frac{M}{\sqrt{2}},$ (36)
that is a Higgs solution. This reduces the set of equations to
$\Box C_{1}(x,x_{1})$ $=$ $-M^{2}C_{1}(x,x_{1})+\delta^{4}(x-x_{1}),$
$\Box C_{2}(x,x_{1},x_{2})$ $=$ $-M^{2}C_{2}(x,x_{1},x_{2})-\frac{6}{\sqrt{2}}MC_{1}(x,x_{2})C_{1}(x,x_{1}),$
$\Box C_{3}(x,x_{1},x_{2},x_{3})$ $=$ $-M^{2}C_{3}(x,x_{1},x_{2},x_{3})-6C_{1}(x,x_{1})C_{1}(x,x_{2})C_{1}(x,x_{3})-\frac{6}{\sqrt{2}}MC_{2}(x,x_{1},x_{2})C_{1}(x,x_{3})$ (37)
$-\frac{6}{\sqrt{2}}MC_{1}(x,x_{1})C_{2}(x,x_{2},x_{3})-\frac{6}{\sqrt{2}}MC_{2}(x,x_{1},x_{3})C_{1}(x,x_{2}),$
$\vdots$ .
For both the solutions, this is signaling the spontaneous breaking of conformal invariance. The Green function is then given by
$C_{1}(p)=\frac{1}{p^{2}+M^{2}},$ (38)
and we have a single boson of mass $M$ with translation invariance never broken.
IV Quantum solution
IV.1 $\Lambda=0$
We can work out the quantum solution exactly in the same way we did for the quartic scalar field in Appendix B. The Dyson-Schwinger equations for our case were obtained in Frasca:2015wva and are the following
$\partial^{2}G_{1}(x)-\mu_{R}^{2}G_{1}(x)+[G_{1}(x)]^{3}+G_{3}(x,x,x)=0,$
$\partial^{2}G_{2}(x,y)-\mu_{R}^{2}G_{2}(x,y)+3[G_{1}(x)]^{2}G_{2}(x,y)=\delta^{4}(x-y),$
$\partial^{2}G_{3}(x,y,z)-\mu_{R}^{2}G_{3}(x,y,z)+3G_{1}^{2}(x)G_{3}(x,y,z)+6G_{1}(x)G_{2}(x,y)G_{2}(x,z)$
$+3G_{2}(x,z)G_{3}(x,x,y)+3G_{2}(x,y)G_{3}(x,x,z)$
$+3G_{2}(x,x)G_{3}(x,y,z)+3G_{1}(x)G_{4}(x,x,y,z)+G_{5}(x,x,x,y,z)=0,$
$\partial^{2}G_{4}(x,y,z,w)-\mu_{R}^{2}G_{4}(x,y,z,w)+3G_{1}^{2}(x)G_{4}(x,y,z,w)+6G_{2}(x,y)G_{2}(x,z)G_{2}(x,w)$
$+6G_{1}(x)G_{2}(x,y)G_{3}(x,z,w)+6G_{1}(x)G_{2}(x,z)G_{3}(x,y,w)+6G_{1}(x)G_{2}(x,w)G_{3}(x,y,z)$
$+3G_{2}(x,y)G_{4}(x,x,z,w)+3G_{2}(x,z)G_{4}(x,x,y,w)$
$+3G_{2}(x,w)G_{4}(x,x,y,z)+3G_{2}(x,x)G_{4}(x,y,z,w)$
$+3G_{1}(x)G_{5}(x,x,y,z,w)+G_{6}(x,x,x,y,z,w)=0,$
$\vdots$ ,
where we have set
$\mu_{R}^{2}=\frac{M^{2}}{2}-3G_{2}(x,x),$ (41)
for the renormalized mass. In the previous section, we have used $C_{1}$ for classical solutions for the Green’s function, which are not correlation functions as intended in quantum field theory. In the case of the quantum solutions, instead of using $C_{1}$, we have used $G_{2}$ to avoid any further confusion in the present context of the computation. In order to get the Gaussian solution out of this set of equations, we note that all the higher order correlation functions ($n>2$) are expressed through $G_{1}$ and $G_{2}$ but, upon integration for the products $G_{2}(x,x^{\prime})G_{2}(x^{\prime},x)$, we observe that these terms have not overlapping support and yields 0222This is a characteristic of distributional products like $\theta(x)\theta(-x)$ where $\theta(x)$ is the Heaviside function.. This has as a consequence that all the correlation functions for $n>2$, when evaluated at equal points, yield 0, largely simplifying the analysis of the equations and producing an exact solution. Thus, we can use the classical solution for $\Lambda=0$ and write down
$G_{1}(x)=\frac{1}{\sqrt{3}}\mu_{R}\text{dn}\left(p\cdot x+\theta,-1\right),$ (42)
where
$p^{2}=\frac{1}{6}\mu_{R}^{2}.$ (43)
This can be seen as a Fubini-Lipatov solution Fubini:1976jm ; Lipatov:1976ny that breaks translation invariance. For the 2P-correlation function, one has
$G_{2}(p)=\frac{\sqrt{2}\pi^{3}}{K^{3}(-1)}\sum_{n=1}^{\infty}n^{2}\frac{e^{-n\pi}}{1+e^{-2n\pi}}\frac{1}{p^{2}+m_{n}^{2}},$ (44)
where $K(-1)$ is the complete elliptic integral of the first kind and
$m_{n}=n\frac{\pi}{K(-1)}\frac{\mu_{R}}{\sqrt{6}}.$ (45)
From the 2P-correlation function we see that the translation invariance is fully recovered in the propagator and the theory displays a discrete spectrum of excitations with a definite mass gap. Anyway, we have a zero mass particle that is not propagating as it does not contribute to the propagator. The mass term gets a quantum correction by
$G_{2}(x,x)=\int\frac{d^{4}p}{(2\pi)^{4}}G_{2}(p),$ (46)
that diverges and must be renormalized. Then, eq.(41) must be solved self-consistently.
For the Higgs-like solution that preserves translation invariance, we observe a mass renormalization given by eq.(41) but modified by
$\mu_{R}^{2}=M^{2}-6G_{2}(x,x).$ (47)
In this case, the quantum effects could wash out the right sign of the mass term modifying the form of the potential that entails the phase transition.
IV.2 $\Lambda\neq 0$
For this case, we consider the Higgs-like solution. We observe that the cubic equation
$\frac{2\sqrt{2}}{3}M\Lambda-\left(-\frac{4}{3}\Lambda+\frac{1}{2}M^{2}\right)\phi+\phi^{3}=0,$ (48)
has 3 real solutions for
$\frac{3}{16}M^{2}>L.$ (49)
The solutions can be written as
$2\sqrt{\frac{1}{3}\left(-\frac{4}{3}\Lambda+\frac{1}{2}M^{2}\right)}\cos\left[\frac{1}{3}\arccos\left[\frac{\sqrt{3}M\Lambda}{2\left(-\frac{4}{3}\Lambda+\frac{1}{2}M^{2}\right)^{\frac{3}{2}}}\right]-\frac{2}{3}\pi k\right],$ (50)
[…]
$\phi_{0}=2\sqrt{\frac{1}{3}\left(-\frac{4}{3}\Lambda+\frac{1}{2}M^{2}\right)}\cos\left[\frac{1}{3}\arccos\left[\frac{\sqrt{3}M\Lambda}{2\left(-\frac{4}{3}\Lambda+\frac{1}{2}M^{2}\right)^{\frac{3}{2}}}\right]-\frac{4}{3}\pi\right].$ (51)
The Dyson-Schwinger set of equations is
$\partial^{2}G_{1}(x)-\mu_{R}^{2}G_{1}(x)+[G_{1}(x)]^{3}+G_{3}(x,x,x)=-\frac{2\sqrt{2}}{3}M\Lambda,$
$\partial^{2}G_{2}(x,y)-\mu_{R}^{2}G_{2}(x,y)+3[G_{1}(x)]^{2}G_{2}(x,y)=\delta^{4}(x-y),$
$\partial^{2}G_{3}(x,y,z)-\mu_{R}^{2}G_{3}(x,y,z)+3G_{1}^{2}(x)G_{3}(x,y,z)+6G_{1}(x)G_{2}(x,y)G_{2}(x,z)$
$+3G_{2}(x,z)G_{3}(x,x,y)+3G_{2}(x,y)G_{3}(x,x,z)$
$+3G_{2}(x,x)G_{3}(x,y,z)+3G_{1}(x)G_{4}(x,x,y,z)+G_{5}(x,x,x,y,z)=0,$
$\partial^{2}G_{4}(x,y,z,w)-\mu_{R}^{2}G_{4}(x,y,z,w)+3G_{1}^{2}(x)G_{4}(x,y,z,w)+6G_{2}(x,y)G_{2}(x,z)G_{2}(x,w)$
$+6G_{1}(x)G_{2}(x,y)G_{3}(x,z,w)+6G_{1}(x)G_{2}(x,z)G_{3}(x,y,w)+6G_{1}(x)G_{2}(x,w)G_{3}(x,y,z)$
$+3G_{2}(x,y)G_{4}(x,x,z,w)+3G_{2}(x,z)G_{4}(x,x,y,w)$
$+3G_{2}(x,w)G_{4}(x,x,y,z)+3G_{2}(x,x)G_{4}(x,y,z,w)$
$+3G_{1}(x)G_{5}(x,x,y,z,w)+G_{6}(x,x,x,y,z,w)=0,$
$\vdots$ ,
where now we have set
$\mu_{R}^{2}=-\frac{4}{3}\Lambda+\frac{M^{2}}{2}-3G_{2}(x,x),$ (54)
that we evaluate in Appendix C. Thus, we have $\phi_{0}$ renormalized by quantum effects. If we write
$M_{R}^{2}=M^{2}-6G_{2}(x,x),$ (55)
we get
$G_{1}=2\sqrt{\frac{1}{3}\left(-\frac{4}{3}\Lambda+\frac{1}{2}M_{R}^{2}\right)}\cos\left[\frac{1}{3}\arccos\left[\frac{\sqrt{3}M_{R}\Lambda}{2\left(-\frac{4}{3}\Lambda+\frac{1}{2}M_{R}^{2}\right)^{\frac{3}{2}}}\right]-\frac{4}{3}\pi\right].$ (56)
This will yield for the 2P-correlation function
$\partial^{2}G_{2}(x,y)+M_{\Lambda}^{2}G_{2}(x,y)=\delta^{4}(x-y).$ (57)
We have introduced the mass of the particle as
$M_{\Lambda}^{2}=\mu_{R}^{2}\left(4\cos^{2}\left[\frac{1}{3}\arccos\left[\frac{\sqrt{3}M_{R}\Lambda}{2\left(-\frac{4}{3}\Lambda+\frac{1}{2}M_{R}^{2}\right)^{\frac{3}{2}}}\right]-\frac{4}{3}\pi\right]-1\right).$ (58)
It reduces to the previous result for $\Lambda=0$ in eq.(47). Differently form the classical model, the quantum one can be solved exactly.
Figure 2: Plot of mass gap.
The scenario here depicted is that of a broken conformal invariance by false vacuum decay through. Correspondingly, a massive particle appears in the theory making the ${\cal R}^{2}$ even more ineffective for the gravitational behavior of the universe. This is true even if the translation invariance is broken by a Fubini-Lipatov instanton for the vacuum as shown for $\Lambda=0$.
In Figure 2, we have plotted the behaviour of the mass gap $M_{\rm\Lambda}$ in a dimensionless unit $M/M_{p}$ with respect to the dimensionless cosmological constant $\Lambda/M^{2}_{p}$. In this plot the blue region is unphysical, as in the regime $\Lambda\gg M^{2}$ which makes the prefactor $\mu^{2}_{R}$ of equation (58) negative. This is clearly visible from equation (54) owing to having an opposite signature in the cosmological constant $\Lambda$ and in the mass gap parameter $M^{2}$, provided $G_{2}(x,x)$ is negligible. On the other hand, in the orange and yellow colored regions, the mass gap $M_{\rm\Lambda}$ increases with the cosmological constant provided the necessary constraint condition $M^{2}\gg\Lambda$ is clearly maintained. In this region due to having this constraint, the prefactor $\mu^{2}_{R}$ of equation (58) is positive. Just like the previous case, when $M^{2}\gg\Lambda$ from equation (54) one can clearly observe that $\mu^{2}_{R}$ is positive. In the right panel of the plot the color coding is helpful in understanding the unphysical as well as the physical region allowed for the present analysis performed in this paper.
V Discussion and Conclusion
  • We have proposed a simple model out of quadratic gravity simplified to the Einstein-Hilbert term plus a quadratic term of the Ricci scalar. This is the Starobinsky model (1). The model is derived assuming that the physics of interest, for strong fields, with conformal symmetry $g_{\mu\nu}=\phi(x)\eta_{\mu\nu}$. Conformal symmetry is broken by quantum effects through the appearance of a mass scale in the quantum theory.
  • The model is well represented through a single mode scalar field characterized by a potential (16) (see Fig. 1) that can display the phenomenon of false vacuum decay, provided a condition is satisfied, already at classical level between the cosmological constant $\Lambda$ and the Starobinsky mass $M$ given by $3M^{2}/16>\Lambda$.
  • The scalar field can be studied and quantized in the limit of a strong field dynamics. For $\Lambda=0$, this yields a mass gap and a spectrum represented in eq.(45). This is a discrete spectrum with an infinite tower of harmonic-oscillator like excitations. The spectrum depends on $G_{2}(x,x)$, a constant to be renormalized. The mass gap scale is proportional to the Starobinsky mass $M$ when $\Lambda$ is set to 0, as can be foreseen.
  • We quantize the model using the technique of Dyson-Schwinger equations described in Appendix B as it is amenable to exact analytical solutions for the 1P- and 2P-correlation functions providing a Gaussian solution for the quantum field theory.
  • The quantized theory provides as a solution for the metric, the 1P-correlation function given in eq.(42), breaking translation invariance being equated to a Fubini-Lipatov instanton and representing the vacuum. This symmetry is recovered at the level of 2P-correlation function (44) that describes the propagating degrees of freedom.
  • When the cosmological constant is different from 0, translation invariance is recovered for $G_{1}$ being a constant and the spectrum (58) (see Fig. 2) is characterized by a single excitation, typical of a Higgs-like behavior. Unphysical regions where there is no propagation exist due to the presence of the cosmological constant.
  • The overall effect of a mass gap in the quadratic gravity makes it negligible at very low-energy where the EH term prevails.
Although we show the development of mass gap in the theory, if that were not the case, that is, we are not in the strong coupling regions and the conditions such as $\Lambda\gg M$ are not satisfied, there would be no way to observe such unusual features: simply because the patches of the Universe where their effects are sizable are not compatible with observers. It is only in the nearly homogeneous and isotropic Hubble patches that cosmic inflation may take place and experimental observations can eventually be done. This is one possible explicit quantum realization of the very popular Linde’s chaotic inflation Linde:1983gd and can be understood at the full quantum level of the classical results of quadratic gravity (for details, see Ref.Salvio:2019ewf ). This may naturally explain the near homogeneity and isotropy of our universe. We plan to study this using the mechanism presented in this paper and also extend this model adding matter content in future work.
Acknowledgement
Authors thank Alexey Koshelev and Alberto Salvio for discussion.
Appendix A: GR computations in generalized harmonic coordinates
In this appendix we show how different choices of gauge with harmonic coordinates can give consistent equations of motion for gravity, assuming a conformally flat solution. This is essential for our argument in the main text.
We have to evaluate the Einstein equations
$g^{\delta\gamma}\partial_{\gamma}\partial_{\delta}g_{\alpha\beta}+\partial_{\beta}g^{\gamma\delta}\partial_{\gamma}g_{\alpha\delta}+\partial_{\alpha}g^{\gamma\delta}\partial_{\gamma}g_{\beta\delta}+2\Gamma^{\gamma}_{\delta\beta}\Gamma^{\delta}_{\gamma\alpha}=0,$ (59)
assuming $g_{\alpha\beta}=\phi\eta_{\alpha\beta}$ the above equation becomes
$\phi^{-1}\Box\phi\eta_{\alpha\beta}-\phi^{-2}\partial_{\beta}\phi\eta^{\gamma\delta}\partial_{\gamma}\phi\eta_{\alpha\delta}-\phi^{-2}\partial_{\alpha}\phi\eta^{\gamma\delta}\partial_{\gamma}\phi\eta_{\beta\delta}+2\Gamma^{\gamma}_{\delta\beta}\Gamma^{\delta}_{\gamma\alpha}$
$=\phi^{-1}\Box\phi\eta_{\alpha\beta}-2\phi^{-2}\partial_{\alpha}\phi\partial_{\beta}\phi+2\Gamma^{\gamma}_{\delta\beta}\Gamma^{\delta}_{\gamma\alpha}=0.$ (60)
Then,
$\Gamma^{\gamma}_{\delta\beta}=\frac{1}{2}g^{\gamma\epsilon}\left[\partial_{\beta}g_{\delta\epsilon}+\partial_{\delta}g_{\beta\epsilon}-\partial_{\epsilon}g_{\delta\beta}\right],$
$\Gamma^{\delta}_{\gamma\alpha}=\frac{1}{2}g^{\delta\rho}\left[\partial_{\alpha}g_{\gamma\rho}+\partial_{\gamma}g_{\alpha\rho}-\partial_{\rho}g_{\gamma\alpha}\right],$ (61)
so that
$\Gamma^{\gamma}_{\delta\beta}=\frac{1}{2}\phi^{-1}\left(\partial_{\beta}\phi\delta^{\gamma}_{\delta}+\partial_{\delta}\phi\delta_{\beta}^{\gamma}-\partial^{\gamma}\phi\eta_{\delta\beta}\right),$
$\Gamma^{\delta}_{\gamma\alpha}=\frac{1}{2}\phi^{-1}\left(\partial_{\alpha}\phi\delta_{\gamma}^{\delta}+\partial_{\gamma}\phi\delta_{\alpha}^{\delta}-\partial^{\delta}\phi\eta_{\gamma\alpha}\right),$ (62)
yielding
$\left(\partial_{\beta}\phi\delta^{\gamma}_{\delta}+\partial_{\delta}\phi\delta_{\beta}^{\gamma}-\partial^{\gamma}\phi\eta_{\delta\beta}\right)\left(\partial_{\alpha}\phi\delta_{\gamma}^{\delta}+\partial_{\gamma}\phi\delta_{\alpha}^{\delta}-\partial^{\delta}\phi\eta_{\gamma\alpha}\right)$
$=4\partial_{\beta}\phi\partial_{\alpha}\phi+\partial_{\beta}\phi\partial_{\alpha}\phi-\partial_{\beta}\phi\partial_{\alpha}\phi+$
$2\partial_{\beta}\phi\partial_{\alpha}\phi-\partial^{\gamma}\phi\partial_{\gamma}\phi\eta_{\alpha\beta}+$
$-\partial_{\beta}\phi\partial_{\alpha}\phi-\partial_{\delta}\phi\partial^{\delta}\phi\eta_{\alpha\beta}+\partial_{\alpha}\phi\partial_{\beta}\phi$
$=6\partial_{\alpha}\phi\partial_{\beta}\phi-2\partial_{\delta}\phi\partial^{\delta}\phi\eta_{\alpha\beta}.$ (63)
We will get
$\phi^{-1}\Box\phi\eta_{\alpha\beta}-2\phi^{-2}\partial_{\beta}\phi\partial_{\alpha}\phi+2\Gamma^{\gamma}_{\delta\beta}\Gamma^{\delta}_{\gamma\alpha}$
$=\phi^{-1}\Box\phi\eta_{\alpha\beta}-2\phi^{-2}\partial_{\beta}\phi\partial_{\alpha}\phi+\phi^{-2}(3\partial_{\alpha}\phi\partial_{\beta}\phi-\partial_{\delta}\phi\partial^{\delta}\phi\eta_{\alpha\beta})$
$\phi^{-1}\Box\phi\eta_{\alpha\beta}+\phi^{-2}\partial_{\beta}\phi\partial_{\alpha}\phi-\phi^{-2}\partial_{\delta}\phi\partial^{\delta}\phi\eta_{\alpha\beta}=0.$ (64)
Now, we introduce
$\phi=e^{f}.$ (65)
This gives
$\Box f\eta_{\alpha\beta}+\partial_{\alpha}f\partial_{\beta}f=0.$ (66)
This equation is not consistent unless $f=constant$. Let us consider the generalized harmonic coordinates then. We will get
$H^{\alpha}=-\phi^{-2}\partial^{\alpha}\phi$ (67)
and then
$H_{\alpha}=-\phi^{-1}\partial_{\alpha}\phi.$ (68)
This gives
$\partial_{\alpha}H_{\beta}+\partial_{\beta}H_{\alpha}=2\phi^{-2}\partial_{\alpha}\phi\partial_{\beta}\phi$ (69)
and
$H_{\delta}\Gamma^{\delta}_{\alpha\beta}=-\phi^{-1}\partial_{\delta}\phi(\frac{1}{2}\phi^{-1}\left(\partial_{\beta}\phi\delta_{\alpha}^{\delta}+\partial_{\alpha}\phi\delta_{\beta}^{\delta}-\partial^{\delta}\phi\eta_{\alpha\beta}\right))=-\phi^{-2}\partial_{\alpha}\phi\partial_{\beta}\phi+\frac{1}{2}\phi^{-2}\partial_{\delta}\phi\partial^{\delta}\phi\eta_{\alpha\beta}.$ (70)
We get
$\phi^{-1}\Box\phi\eta_{\alpha\beta}+\phi^{-2}\partial_{\beta}\phi\partial_{\alpha}\phi+2\phi^{-2}\partial_{\alpha}\phi\partial_{\beta}\phi+2\phi^{-2}\partial_{\alpha}\phi\partial_{\beta}\phi-2\phi^{-2}\partial_{\delta}\phi\partial^{\delta}\phi\eta_{\alpha\beta}=0.$ (71)
This equation is inconsistent yet. The only way around is to consider the function $H_{\alpha}$ arbitrary as in in lattice computations Pretorius:2006tp to remove the offending terms. We will have
$\phi^{-1}\Box\phi\eta_{\alpha\beta}+\phi^{-2}\partial_{\beta}\phi\partial_{\alpha}\phi-\phi^{-2}\partial_{\delta}\phi\partial^{\delta}\phi\eta_{\alpha\beta}+(\partial_{\alpha}H_{\beta}+\partial_{\beta}H_{\alpha})-2H_{\delta}\Gamma^{\delta}_{\alpha\beta}=0.$ (72)
For consistency reasons we have to chose
$-\phi^{-2}\partial_{\beta}\phi\partial_{\alpha}\phi=(\partial_{\alpha}H_{\beta}+\partial_{\beta}H_{\alpha})+\phi^{-1}\left(\partial_{\beta}\phi H_{\alpha}+\partial_{\alpha}\phi H_{\beta}-H_{\delta}\partial^{\delta}\phi\eta_{\alpha\beta}\right).$ (73)
The choice $\chi/M_{p}=\log\phi$ fixes the gauge properly.
Appendix B: Derivation of Dyson-Schwinger equations for a quartic scalar field
The technique we use to derive the set of Dyson-Schwinger equations in PDE shape has been devised by Bender, Milton and Savage Bender:1999ek for a PT-invariant quantum mechanical model. We extended this approach to quantum field theory Frasca:2015yva . For our convenience, we work with an Euclidean metric and write the partition function as
$Z[j]=\int[d\phi]e^{-\int d^{4}x[\frac{1}{2}(\partial\phi)^{2}-\frac{\lambda}{4}\phi^{4}+j\phi]}$ (74)
where $j$ is arbitrary source and $\lambda$ the theory self-coupling. Our aim is to evaluate the full set of nP-correlation functions
$G_{n}(x_{1},x_{2},\ldots,x_{n})=\left.\frac{\delta^{n}\ln Z}{\delta j(x_{1})\delta j(x_{2})\ldots\delta j(x_{n})}\right|_{j=0},$ (75)
so that we are able to get a complete solution for the quantum field theory and all the observables like decay rates and scattering cross sections will be computable through the LSZ theorem. It is straightforward to write down
$G_{1}(x_{1})$ $=$ $\left.\frac{1}{Z}\frac{\delta Z}{\delta j(x_{1})}\right|_{j=0},$
$G_{2}(x_{1},x_{2})$ $=$ $\left.\frac{\delta G_{1}^{(j)}(x_{1})}{\delta j(x_{2})}\right|_{j=0},$
$G_{3}(x_{1},x_{2},x_{3})$ $=$ $\left.\frac{\delta G_{2}^{(j)}(x_{1},x_{2})}{\delta j(x_{3})}\right|_{j=0},$ (76)
$\vdots$ ,
where
$G_{n}^{(j)}(x_{1},x_{2},\ldots,x_{n})=\frac{\delta^{n}\ln Z}{\delta j(x_{1})\delta j(x_{2})\ldots\delta j(x_{n})},$ (77)
that are the correlation functions dependent on the source $j$. The equation of motion takes the form
$\partial^{2}\phi+\lambda\phi^{3}=j.$ (78)
In order to get $G_{1}$, we average this equation with $Z$ obtaining
$\partial^{2}G_{1}^{(j)}(x)+\lambda Z^{-1}\langle\phi^{3}\rangle=j,$ (79)
but using
$ZG_{1}^{(j)}(x)=\langle\phi\rangle,$ (80)
we can iterate the derivative with respect to $j(x)$ obtaining
$ZG_{2}^{(j)}(x,x)+Z[G_{1}^{(j)}(x)]^{2}=\langle\phi^{2}\rangle,$
$3ZG_{1}^{(j)}(x)G_{2}^{(j)}(x,x)+ZG_{3}^{(j)}(x,x,x)+Z[G_{1}^{(j)}(x)]^{3}=\langle\phi^{3}\rangle.$ (81)
By direct substitution in eq.(79), we get the PDE for $G_{1}^{(j)}$
$\partial^{2}G_{1}^{(j)}(x)+\lambda\left\{3G_{2}^{(j)}(x,x)G_{1}^{(j)}(x)+G_{3}^{(j)}(x,x,x)+[G_{1}^{(j)}(x)]^{3}\right\}=j,$ (82)
and, by setting $j=0$, one has
$\partial^{2}G_{1}(x)+\lambda[3G_{1}(x)G_{2}(x,x)+G_{3}(x,x,x)+G_{1}^{3}(x)]=0.$ (83)
To get the equation for $G_{2}$. we need to derive eq.(82) with respect to $j(x_{2})$. This will yield
$\partial^{2}G_{2}^{(j)}(x,x_{2})+\lambda\left\{3G_{2}^{(j)}(x,x)G_{2}^{(j)}(x,x_{2})+3G_{2}^{(j)}(x,x)G_{3}^{(j)}(x,x,x_{2})+G_{4}^{(j)}(x,x,x,x_{2})+3[G_{1}^{(j)}(x)]^{2}G_{2}^{(j)}(x,x_{2})\right\}.$ (84)
By setting $j=0$, one gets the equation for $G_{2}$. The procedure can be iterated at any order to compute the higher-order correlation functions. This approach is successful when exact solution are known for $G_{1}$ and $G_{2}$. Only in this case, a Gaussian solution for the partition function can be obtained by selecting $G_{n}=0$ for $n>2$, provided at least two independent variables are identical.
Appendix C: Mass shift
We want to evaluate the contribution of the mass shift in eq.(54) by evaluating the self-consistency equation
$\mu_{R}^{2}=-\frac{4}{3}\Lambda+\frac{M^{2}}{2}-3\int\frac{d^{4}p}{(2\pi)^{4}}\frac{1}{p^{2}+\mu_{R}^{2}},$ (85)
moving to the Euclidean metric in the integral. The integral can be regularized by a cut-off $K$ yielding
$\mu_{R}^{2}=-\frac{4}{3}\Lambda+\frac{M^{2}}{2}-\frac{3}{16\pi^{2}}\left[K^{2}-\mu_{R}^{2}\log\left(1+\frac{K^{2}}{\mu_{R}^{2}}\right)\right].$ (86)
This can be solved by iteration starting with $\mu_{R}=0$ and noticing that, for $M$ very large, the theory has a natural cut-off for $K=M$. Thus, while this correction can appear to be small, it could not be so for the cosmological constant term.