### v2

# Mass gap in non-perturbative quadratic $\mathcal{R}^{2}$ gravity via Dyson-Schwinger

Abstract
We apply in a simple model derived from quadratic $\mathcal{R}^{2}$ gravity the technique of Dyson-Schwinger equations to solve for its corresponding quantum theory. Particularly, we solve the classical equations of motion to get a solution to the hierarchy of Dyson-Schwinger equations in the limit of large Ricci scalar, assumed to be constant and larger than the square of the Starobinsky mass. Moving to the Einstein frame, the model admits Higgs-like solutions with a single particle having a finite mass. We quantize the scalar field showing the appearing of a mass gap through a Higgs-like solution. The presence of the mass gap, that increases with the square root of the Ricci scalar, shows how the effect of the scalar sector at low-energy becomes ineffective, making it relevant only at short distances.
I Introduction
One of the most successful theories of the past century and certainly the most successful theory in the context of gravity is Albert Einstein’s general relativity (GR). This theory has stood the test of time again and again which has been verified by several observations. These experimental observations range from those as recent as the discovery of gravitational waves, with coalescing binary black hole astrophysical systems LIGOScientific:2016aoc , to the extraordinary optical images of the black holes in the center of our galaxy and the M87 one as captured by the Event Horizon Telescope EventHorizonTelescope:2019dse ; EventHorizonTelescope:2019ggy ; EventHorizonTelescope:2022xnr ; EventHorizonTelescope:2022wok ; EventHorizonTelescope:2022xqj . Extending GR, the theory suitably can account for matter fields, and forms the edifice of our present understanding of modern day cosmology, including the early universe right from the origin of hot big bang.
Even with so much success at large scales, nonetheless, at the quantum frontier GR is known to be non-renormalizable by perturbative methods Goroff:1985sz ; Goroff:1985th and may not remain perturbative at length scales much below the Planck one. Also it encounters Big Bang and Black Hole singularity problems and known to pathological behaviour in the ultra-violate (UV) Salvio:2018crh . GR also suffers from the what is known as the “conformal-factor problem”: in simple terms this is nothing but absence of a consistent Euclidean path integral. Moreover this becomes quite serious since all known quantum field theories (QFT) can be consistently defined only as analytic continuations of the corresponding Euclidean version, see Ref. Salvio:2024joi for details.
The issue of non-renormalizability and the issue of conformal-factor is solved when GR is augmented with quadratic-in-curvature terms which is often called quadratic gravity (See Salvio:2018crh ; Donoghue:2021cza for reviews.). It has been shown to be renormalizable Weinberg:1974tw ; Deser:1975nv ; Stelle:1976gc ; Barvinsky:2017zlx and following the prescription of Ref. Gibbons:1994cg one may determine the Euclidean action which implies that it is also free from the conformal-factor problem in most of the regions of its parameter space Menotti:1989ms . Such quadratic gravity theories also address the singularity issues. The classical Hamiltonian of quadratic gravity is unbounded from below, something known in the literature as the Ostrogradsky’s instability issue as shown in Ref. ostro . Frankly, this is common in all kinds of higher-derivative gravity extensions. Nonetheless estimations of runaway rates show that such a decay can be slow enough to be compatible with observations as studied in Refs. Salvio:2019ewf ; Gross:2020tph ; Held:2021pht ; Held:2023aap .
In this paper we offer a non-perturbative formulation of quadratic $\mathcal{R}^{2}$ gravity. This is based on the Dyson-Schwinger approach and may certainly provide us with a fully quantum and non-perturbative justification for these classical cosmological findings in the long run. Furthermore, such a formulation that we show here can also be utilised to validate and check consistency with other independent perturbative methods and sometimes even background-dependent approaches to formulate quadratic gravity which been shown to preserve unitarity and (meta)stablity. In the context of renormalizable theories, these approaches include the Lee-Wick Lee:1969fy ; Salvio:2018kwh ; Donoghue:2019fcb ; Holdom:2021hlo the fakeon prescription Anselmi:2017ygm and string-inspired non-local ghost-free approaches to higher-derivative gravity Frasca:2020jbe ; Frasca:2020ojd ; Frasca:2021iip ; Frasca:2022duz ; Frasca:2022gdz . Particularly for quadratic gravity this has been studied rigorously in Ref. Salvio:2024joi .
The novel tool to treat the non-perturbative systems in the framework of Dyson-Schwinger Equations that we propose in this paper are based on the exact solution of the background equations of motion in terms of Jacobi elliptic functions. This is inspired by the analytic approach of Dyson-Schwinger equations (DSE), originally devised by Bender, Milton and Savage Bender:1999ek . Using this we are able to represent the Green’s functions of the theory purely analytically, therefore we understand the effect of the background on the interactions that remains to be valid even in the strong coupling regimes Frasca:2015yva . This approach has been widely used to investigate QCD with great success with respect to results obtain closely matching with lattice QCD results Frasca:2021yuu ; Frasca:2021mhi ; Frasca:2022lwp ; Frasca:2022pjf ; Chaichian:2018cyv and to the scalar sector Frasca:2015wva  [^1] Moreover extensions beyond the traditional QFT and SM of particle physics to other types of models involving new gauge sectors and string-inspired non-local theories have been widely discussed using this methodology Frasca:2019ysi ; Chaichian:2018cyv ; Frasca:2017slg ; Frasca:2016sky ; Frasca:2015yva ; Frasca:2015wva ; Frasca:2013tma ; Frasca:2012ne ; Frasca:2009bc ; Frasca:2010ce ; Frasca:2008tg ; Frasca:2009yp ; Frasca:2008zp ; Frasca:2007uz ; Frasca:2006yx ; Frasca:2005sx ; Frasca:2005mv ; Frasca:2005fs . In context to experiments concerning particle physics phenomenology like non-perturbative hadronic vacuum polarization contributions to the muon anomalous magnetic moment $(g-2)_{\mu}$ was predicted in Ref.Frasca:2021yuu , QCD in the non-perturbative regimes Frasca:2021mhi ; Frasca:2022lwp ; Frasca:2022pjf , Higgs-Yukawa theory Frasca:2023qii , finite temperature field theory Frasca:2023eoj , non-perturbative false vacuum transitions leading to phase transitions Frasca:2022kfy ; Calcagni:2022tls ; Calcagni:2022gac , dark energy Frasca:2022vvp , and investigations of the mass gap and confinement in string-inspired infinite-derivative and Lee-Wick theories Frasca:2020jbe ; Frasca:2020ojd ; Frasca:2021iip ; Frasca:2022duz ; Frasca:2022gdz . These tools were also employed to implement dynamical generation electroweak (EW) scale via scalar, vector boson and fermionic condensates too very recently Frasca:2024fuq ; Chatterjee:2024dgw ; Frasca:2024pmv ; Frasca:2023eoj ; Frasca:2024ame as an approach to address the gauge hierarchy problem in the Standard Model (SM) of particle physics.
A final formulation of the theory should involve quantum cosmology with a quantum mechanical “wave function of the universe”, as investigated by Hawking and Hartle Hartle:1983ai . This actually can be used to solve the Schrodinger equation, also known as Wheeler-DeWitt equation in this context DeWitt:1967yk ; Wheeler . Therefore a fully consistent quantum cosmology with GR extended to quadratic gravity is what we allude to, since this theory contains Starobinsky’s model of inflation (also known as $\mathcal{R}^{2}$ inflation) Starobinsky:1980te , which is currently in perfect agreement with the cosmic microwave background radiation (CMBR) observations assuming the cosmic inflationary paradigm Ade:2015lrj ; Planck2018:inflation ; BICEP:2021xfz .
Yet another application of a non-perturbative quadratic gravity via Dyson-Schwinger approach would be in the paradigm of asymptotic safety originally proposed by Steve Weinberg WeinbergAS , where gravity is made ultra-violate (UV) complete and shown to be renormalizable due to presence of a non-perturbative UV fixed point. Investigations of such fixed points have been extensively carried out for quadratic $\mathcal{R}^{2}$ gravity in Refs. Benedetti:2009rx ; Falls:2020qhj which lead to consistent realization of Asymptotic Safe Gravity.
As we will show that the Einstein-Hilbert term plus a quadratic term of the Ricci scalar, for strong fields. Conformal symmetry is broken by the appearance of a mass gap in the theory. The mass spectrum of theory is a discrete spectrum with an infinite tower of harmonic-oscillator like excitations. The mass gap scale is proportional to the Starobinsky mass which is the coefficient of the $\mathcal{R}^{2}$ term which represents the self-interactions among the gravitons and the scalar degrees of freedom.
In this work, we start considering generally non-renormalizable theories and obtain some meaningful renormalizable scalar theories to work with, only under some conditions. There exists several such examples, a well-known example being given by the Nambu-Jona-Lasinio model used to describe strong interactions Klevansky:1992qe . This model is characterized by fermion fields with a quartic interaction between them. As such, this model is not renormalizable and the computations depend on a cut-off. When a bosonization procedure is applied Klevansky:1992qe , a $\sigma$-model emerges that is renormalizable, and could represent a description of the breaking of the chiral symmetry in strong interactions. With this principle in mind, we try to have similar approach here showing that, under proper conditions, some effective scalar models could be extracted from quadratic gravity and can be suitably properly studied using the well-known techniques of quantum field theory as the Dyson-Schwinger equations are. However a concrete re-normalizable model of gravity like Stelle gravity or asymptotic safe gravity or non-local gravity (see Ref. Salvio:2018crh for such a discussion) is beyond the scope of the current paper and will be studied in future.
The paper is organized as follows: In Sec. II, we introduce the model and derive the scalar field theory arising from it. In Sec. III, we discuss the classical solutions for the scalar field theory. In Sec. IV, we quantize the model ans solve the set of Dyson-Schwinger equations both with a null or non-null the cosmological constant. In Sec.V, we provide the conclusions.
II Scalar Model for Quadratic Starobinsky gravity: A small review
We present a scalar field model from the full action of the quadratic gravity Starobinsky:1980te ; Mukhanov:1981xt ; Kaneda:2010ut and then we will quantize it with the technique of the Dyson-Schwinger equations [^2] We do not consider here a scenario like Higgs inflation, involving a term like $\xi\phi^{2}\mathcal{R}$. We only restrict ourselves to a minimally coupled scalar field to gravity as comes out from quadratic gravity..
$S=-\frac{M_{p}^{2}}{2}\int d^{4}x\sqrt{-g}\left(\mathcal{R}-2\Lambda+\frac{\mathcal{R}^{2}}{6M^{2}}\right),$ (1)
where $M_{p}$ is the Planck mass and $M$ is a new mass scale in this theory which is called the Starobinsky mass or also known as the conformal coupling. This action can be re-written in terms of the interaction between gravity and a scalar field $\chi$ after performing the transformation Whitt:1984pd ; Kehagias:2013mya
$g_{\mu\nu}\rightarrow e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}g_{\mu\nu},$ (2)
where $g_{\mu\nu}$ is a generic metric. This gives Salvio:2018crh :
$S=\int{\rm d}^{4}x\sqrt{-g}\left[-\frac{M_{\rm p}^{2}}{2}(\mathcal{R}-2\Lambda e^{-2\sqrt{\frac{2}{3}}\frac{\chi}{M_{p}}})+\frac{1}{2}g^{\mu\nu}\partial_{\mu}\chi\partial_{\nu}\chi+\frac{3}{4}M_{\rm p}^{2}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}\right)^{2}\right].$ (3)
This action yields the following sets of equations of motion, which are given by $\frac{\delta S}{\delta g_{\mu\nu}}$ and $\frac{\delta S}{\delta\chi}$:
$\mathcal{R}_{\mu\nu}=\frac{1}{M_{p}^{2}}\left\{\partial_{\mu}\chi\,\partial_{\nu}\chi-\tfrac{1}{2}g_{\mu\nu}\,g^{\alpha\beta}\partial_{\alpha}\chi\,\partial_{\beta}\chi-g_{\mu\nu}\left[M_{\rm p}^{2}\Lambda\,e^{-2\sqrt{\tfrac{2}{3}}\tfrac{\chi}{M_{\rm p}}}+\tfrac{3}{4}M_{\rm p}^{2}M^{2}\left(1-e^{-\sqrt{\tfrac{2}{3}}\tfrac{\chi}{M_{\rm p}}}\right)^{2}\right]\right\}+\frac{1}{2}g_{\mu\nu}{\cal R},$
$\Box_{g}\chi=\sqrt{\frac{3}{2}}M^{2}M_{p}e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{p}}}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{p}}}\right)-\sqrt{\frac{8}{3}}\Lambda M_{p}e^{-2\sqrt{\frac{2}{3}}\frac{\chi}{M_{p}}}$ (4)
where $\Box_{g}=-\frac{1}{\sqrt{-g}}\partial_{\mu}g^{\mu\nu}\sqrt{-g}\partial_{\nu}$ is the Laplace-Beltrami operator.
We look for solution with $R$ being a constant. This yield the set of equations for the scalar field
$\partial_{\sigma}\chi\partial^{\sigma}\chi+3M_{\rm p}^{2}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}\right)^{2}+4M_{p}^{2}\Lambda e^{-2\sqrt{\frac{2}{3}}\frac{\chi}{M_{p}}}-M_{p}^{2}\mathcal{R}=0,$
$-\frac{1}{\sqrt{-g}}\partial_{\mu}g^{\mu\nu}\sqrt{-g}\partial_{\nu}\chi=\sqrt{\frac{3}{2}}M_{\rm p}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}\right)e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}-\sqrt{\frac{8}{3}}\Lambda M_{p}e^{-2\sqrt{\frac{2}{3}}\frac{\chi}{M_{p}}}.$ (5)
Taking a dimensionless rescaled field variable $\varphi=\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}$, we further get:
$\partial_{\sigma}\varphi\partial^{\sigma}\varphi+2M^{2}\left(1-e^{-\varphi}\right)^{2}+\frac{8}{3}\Lambda e^{-2\varphi}-\frac{2}{3}\mathcal{R}=0$
$\Box_{g}\varphi=M^{2}\left(1-e^{-\varphi}\right)e^{-\varphi}-\frac{4}{3}\Lambda e^{-2\varphi}.$ (6)
Let us take $\varphi=-\log Z$, we get
$Z^{-2}\partial_{\sigma}Z\partial^{\sigma}Z+2M^{2}\left(1-Z\right)^{2}+\frac{8}{3}\Lambda Z^{2}-\frac{2}{3}\mathcal{R}=0$
$-Z^{-1}\Box_{g}Z+Z^{-2}\partial_{\sigma}Z\partial^{\sigma}Z=M^{2}\left(1-Z\right)Z-\frac{4}{3}\Lambda Z^{2}.$ (7)
We insert the first equation into the second obtaining
$\Box_{g}Z=-\left(2M^{2}-\frac{2}{3}\mathcal{R}\right)Z+3M^{2}Z^{2}-\left(M^{2}+\frac{4}{3}\Lambda\right)Z^{3}.$ (8)
Thus, we are left with a quartic scalar $\phi^{4}$-like theory which can be easily seen upon the introduction of the following mapping of variables:
$Z=\frac{M^{2}}{M^{2}+\frac{4}{3}\Lambda}+\frac{\phi}{\sqrt{M^{2}+\frac{4}{3}\Lambda}},$ (9)
that gives us
$\Box_{g}\phi=-\left(2M^{2}-\frac{2}{3}\mathcal{R}\right)\frac{M^{2}}{\sqrt{M^{2}+\frac{4}{3}\Lambda}}+\frac{2M^{6}}{\left(M^{2}+\frac{4}{3}\Lambda\right)^{\frac{3}{2}}}$ (10)
$+\left[-\left(2M^{2}-\frac{2}{3}\mathcal{R}\right)+\frac{3M^{4}}{M^{2}+\frac{4}{3}\Lambda}\right]\phi-\phi^{3}$
Such a potential could signal a false vacuum decay yielding, after integrating on $\phi$ and changed sign to the rhs of eq.(10),
$V(\phi)=\left[\left(2M^{2}-\frac{2}{3}\mathcal{R}\right)\frac{M^{2}}{\sqrt{M^{2}+\frac{4}{3}\Lambda}}-\frac{2M^{6}}{\left(M^{2}+\frac{4}{3}\Lambda\right)^{\frac{3}{2}}}\right]\phi-\frac{1}{2}\left[-\left(2M^{2}-\frac{2}{3}\mathcal{R}\right)+\frac{3M^{4}}{M^{2}+\frac{4}{3}\Lambda}\right]\phi^{2}+\frac{1}{4}\phi^{4}.$ (11)
We have set the integration constant to 0.
Figure 1: Plot of the potential of the conformal factor. It is seen the possibility of false vacuum decay for $3M^{2}<\mathcal{R}$. All the physical quantity are dimensionless being scaled with the Planck mass $M_{P}$.
Fig. 1 shows different potentials varying the Starobinsky mass $M$ with respect to the Planck mass for a given value of the cosmological constant. The characteristic shape of a false vacuum decay potential can be recognized provided the condition $3M^{2}<\mathcal{R}$ holds.
The condition $3M^{2}<\mathcal{R}$ (strong coupling regime) is essential for our analysis. It grants the reduction of the Laplace-Beltrami operator to that of the flat space, aside from a proper scale changing. Physically, this means that the quadratic term $\mathcal{R}^{2}$ is prevailing with respect to the linear one of the standard Einstein-Hilbert action. For this reason, the quantum behavior in this regime is the main concern of the present study.
We see that the presence of the cosmological constant is not so relevant and so, in the following we set it to 0. This will yield the equation of motion
$\Box_{g}\phi=\frac{2}{3}\mathcal{R}M+\left[\frac{2}{3}\mathcal{R}+M^{2}\right]\phi-\phi^{3},$ (12)
and the potential
$V(\phi)=-\frac{2}{3}\mathcal{R}M\phi-\frac{1}{2}\left[\frac{2}{3}\mathcal{R}+M^{2}\right]\phi^{2}+\frac{\phi^{4}}{4}.$ (13)
III Classical solutions
For our study, we need to introduce an arbitrary source $j$ and compute the functional series
$\phi(x)=\phi_{0}(x)+\sum_{n=1}^{\infty}\int C_{n}(x,x_{1},\ldots,x_{n-1})\prod_{k=1}^{n}j(x_{k})d^{4}x_{k},$ (14)
where we are assuming $\phi$ a functional of an arbitrary source $j$ and
$C_{n}(x,x_{1},\ldots,x_{n-1})=\left.\frac{\delta^{n}\phi(x)}{\delta j(x_{1})\ldots\delta j(x_{n-1})}\right|_{j=0}.$ (15)
By iterating the derivation with respect to $j$ and starting from the equation of motion (that holds for $3M^{2}<\mathcal{R}$),
$\Box\phi(x)=\Omega+\mu_{0}^{2}\phi(x)-\phi^{3}(x)+j(x),$ (16)
where $\Omega=\frac{2}{3}\mathcal{R}M$ and $\mu_{0}^{2}=\frac{2}{3}\mathcal{R}+M^{2}=\frac{\Omega}{M}+M^{2}$ having set $\Lambda=0$ to make handling the equations simpler, we get
$\Box\phi_{0}(x)$ $=$ $\Omega+\mu_{0}^{2}\phi_{0}(x)-\phi_{0}^{3}(x),$
$\Box C_{1}(x,x_{1})$ $=$ $\left[\mu_{0}^{2}-3\phi_{0}^{2}(x)\right]C_{1}(x,x_{1})+\delta^{4}(x-x_{1}),$
$\Box C_{2}(x,x_{1},x_{2})$ $=$ $\left[\mu_{0}^{2}-3\phi_{0}^{2}(x)\right]C_{2}(x,x_{1},x_{2})-6\phi_{0}(x)C_{1}(x,x_{2})C_{1}(x,x_{1}),$
$\Box C_{3}(x,x_{1},x_{2},x_{3})$ $=$ $\left[\mu_{0}^{2}-3\phi_{0}^{2}(x)\right]C_{3}(x,x_{1},x_{2},x_{3})-6C_{1}(x,x_{1})C_{1}(x,x_{2})C_{1}(x,x_{3})-6\phi_{0}(x)C_{2}(x,x_{1},x_{2})C_{1}(x,x_{3})$ (17)
$-6\phi_{0}(x)C_{1}(x,x_{1})C_{2}(x,x_{2},x_{3})-6\phi_{0}(x)C_{2}(x,x_{1},x_{3})C_{1}(x,x_{2}),$
$\vdots$ .
The cubic polynomial on the lhs of the leading order equation has three solution as expected $(-M,(3M-\sqrt{3}\sqrt{3M^{2}+8R})/6,(3M+\sqrt{3}\sqrt{3M^{2}+8R})/6)$. The last one is the true vacuum. Thus, if we select it, we can work out a Higgs-like solution and a massive solution already at this stage. We will discuss this point below. On the other hand, we can exploit the strong coupling condition $3M^{2}<\mathcal{R}$ and neglect the constant $\Omega$ with respect to the other terms in the leading order equation. We can also approximate $\mu_{0}^{2}\approx 2R/3$ that can be seen as a leading order approximation [^3] We observe that we are claiming that $RM\ll R\phi_{0}$ that is $\phi_{0}\gg M$ that is consistent with our approximation discussed in the previous section.. This implies that the leading order equation takes the form
$\Box\phi_{0}(x)\approx\mu_{0}^{2}\phi_{0}(x)-\phi_{0}^{3}(x),$ (18)
being $\mu_{0}^{2}$ the approximate value given above. In this way, we are able to get an alternative approximate solution to this set of equations yielding
$\phi_{0}(x)=\sqrt{\frac{2}{3}}\mu_{0}\text{dn}\left(p\cdot x+\theta,-1\right),$ (19)
where one can see M as the Starobinsky mass in the original Lagrangian (1), dn is a Jacobian elliptic function and
$p^{2}=\frac{\mu_{0}^{2}}{3}.$ (20)
Similarly, also in this case
$C_{1}(p)=\frac{\sqrt{2}\pi^{3}}{K^{3}(-1)}\sum_{n=1}^{\infty}n^{2}\frac{e^{-n\pi}}{1+e^{-2n\pi}}\frac{1}{p^{2}+m_{n}^{2}},$ (21)
where K(1)K(-1) is the complete elliptic integral of the first kind and
$m_{n}=n\frac{\pi}{K(-1)}\frac{\mu_{0}}{\sqrt{3}}.$ (22)
We can iterate obtaining
$C_{2}(x,x_{1},x_{2})$ $=$ $-6\int d^{4}x^{\prime}C_{1}(x,x^{\prime})\phi_{0}(x^{\prime})C_{1}(x^{\prime},x_{2})C_{1}(x^{\prime},x_{1}),$
$C_{3}(x,x_{1},x_{2},x_{3})$ $=$ $-6\int d^{4}x^{\prime}C_{1}(x,x^{\prime})C_{1}(x^{\prime},x_{1})C_{1}(x^{\prime},x_{2})C_{1}(x^{\prime},x_{3})-6\int d^{4}x^{\prime}C_{1}(x,x^{\prime})\phi_{0}(x^{\prime})C_{2}(x^{\prime},x_{1},x_{2})C_{1}(x^{\prime},x_{3})$ (23)
$-6\int d^{4}x^{\prime}C_{1}(x,x^{\prime})\phi_{0}(x^{\prime})C_{1}(x^{\prime},x_{1})C_{2}(x^{\prime},x_{2},x_{3})-6\int d^{4}x^{\prime}C_{1}(x,x^{\prime})\phi_{0}(x^{\prime})C_{2}(x^{\prime},x_{1},x_{3})C_{1}(x^{\prime},x_{2}),$
$\vdots$ .
We see that our solution is completely obtained by knowing $\phi_{0}$ and the Green function $C_{1}(x,x^{\prime})$ that is translation invariant and can be written as $C_{1}(x-x^{\prime})$. We will see that the same result can be obtained in quantum field theory providing a Gaussian solution to the set of Dyson-Schwinger equations of the theory.
The Higgs-like case can analyzed by taking
$\phi_{0}\approx\mu_{0}.$ (24)
This reduces the set of equations to
$\Box C_{1}(x,x_{1})$ $=$ $-2\mu_{0}^{2}C_{1}(x,x_{1})+\delta^{4}(x-x_{1}),$
$\Box C_{2}(x,x_{1},x_{2})$ $=$ $-2\mu_{0}^{2}C_{2}(x,x_{1},x_{2})-\frac{6}{\sqrt{2}}MC_{1}(x,x_{2})C_{1}(x,x_{1}),$
$\Box C_{3}(x,x_{1},x_{2},x_{3})$ $=$ $-2\mu_{0}^{2}C_{3}(x,x_{1},x_{2},x_{3})-6C_{1}(x,x_{1})C_{1}(x,x_{2})C_{1}(x,x_{3})-\frac{6}{\sqrt{2}}MC_{2}(x,x_{1},x_{2})C_{1}(x,x_{3})$ (25)
$-\frac{6}{\sqrt{2}}MC_{1}(x,x_{1})C_{2}(x,x_{2},x_{3})-\frac{6}{\sqrt{2}}MC_{2}(x,x_{1},x_{3})C_{1}(x,x_{2}),$
$\vdots$ .
For both the solutions, this is signaling the spontaneous breaking of conformal invariance. The Green function is then given by
$C_{1}(p)=\frac{1}{p^{2}+2\mu_{0}^{2}},$ (26)
and we have a single boson of mass $\sqrt{2}\mu_{0}$ with translation invariance never broken after quantization. We see that the classical theory displays a massive solution in any case, due to the nonlinear potential we obtained from the Starobinsky model.
IV Quantum solution
From the analysis in the precedent section, the true minimum is located at
$G_{1}=\phi_{0}=\frac{1}{6}(3M+\sqrt{3}\sqrt{3M^{2}+8R})\approx\sqrt{\frac{2}{3}R}=\mu_{0},$ (27)
in the considered approximation $3M^{2}<\mathcal{R}$, given the Dyson-Schwinger set of equations
$\partial^{2}G_{1}(x)-\mu_{R}^{2}G_{1}(x)+[G_{1}(x)]^{3}+G_{3}(x,x,x)=\Omega,$
$\partial^{2}G_{2}(x,y)-\mu_{R}^{2}G_{2}(x,y)+3[G_{1}(x)]^{2}G_{2}(x,y)=\delta^{4}(x-y),$
$\partial^{2}G_{3}(x,y,z)-\mu_{R}^{2}G_{3}(x,y,z)+3G_{1}^{2}(x)G_{3}(x,y,z)+6G_{1}(x)G_{2}(x,y)G_{2}(x,z)$
$+3G_{2}(x,z)G_{3}(x,x,y)+3G_{2}(x,y)G_{3}(x,x,z)$
$+3G_{2}(x,x)G_{3}(x,y,z)+3G_{1}(x)G_{4}(x,x,y,z)+G_{5}(x,x,x,y,z)=0,$
$\partial^{2}G_{4}(x,y,z,w)-\mu_{R}^{2}G_{4}(x,y,z,w)+3G_{1}^{2}(x)G_{4}(x,y,z,w)+6G_{2}(x,y)G_{2}(x,z)G_{2}(x,w)$
$+6G_{1}(x)G_{2}(x,y)G_{3}(x,z,w)+6G_{1}(x)G_{2}(x,z)G_{3}(x,y,w)+6G_{1}(x)G_{2}(x,w)G_{3}(x,y,z)$
$+3G_{2}(x,y)G_{4}(x,x,z,w)+3G_{2}(x,z)G_{4}(x,x,y,w)$
$+3G_{2}(x,w)G_{4}(x,x,y,z)+3G_{2}(x,x)G_{4}(x,y,z,w)$
$+3G_{1}(x)G_{5}(x,x,y,z,w)+G_{6}(x,x,x,y,z,w)=0,$
$\vdots$ ,
where now we have set
$\mu_{R}^{2}=\mu_{0}^{2}-3G_{2}(x,x),$ (30)
with the correction due to quantum fluctuations evaluated in Appendix B. Thus, we have $\phi_{0}$ renormalized by quantum effects. This will yield for the 2P-correlation function
$\partial^{2}G_{2}(x,y)+M_{G}^{2}G_{2}(x,y)=\delta^{4}(x-y),$ (31)
where the final Mass Gap developed in the theory reads as (see Appendix B)
$M^{2}_{G}=2\mu_{R}^{2}=\frac{8\pi^{2}\left(\sqrt{3}\sqrt{3M^{2}+8R}+3M\right)^{2}}{27W\left(\frac{16}{27}\pi^{3}\left(\sqrt{3}\sqrt{3M^{2}+8R}+3M\right)^{2}e^{-1+\gamma+\frac{16\pi^{2}}{3}}\right)},$ (32)
where $W(x)$ is the Lambert function that solves the equation $xe^{x}=w$.
Figure 2: Plot of mass gap.
We observe that the conformal factor that appears in passing from the Jordan to the Einstein frame, yields a propagating scalar field. The appearance of a mass gap in the quantum study this Starobinsky scalar field displays a breaking of the scale symmetry, reducing the geometry of the spacetime to that of the backbone geometry that has a null Ricci curvature. This shows a transition between a full blown quadratic gravity to a Einstein-Hilbert dominated gravity.
[…]
V Discussion and Conclusion
  • We have proposed a simple model out of quadratic gravity simplified to the Einstein-Hilbert term plus a quadratic term of the Ricci scalar. This is the Starobinsky model (1). The model is derived assuming that the physics of interest, for strong fields, is described by a large and constant Ricci scalar granting the appearance of a mass gap in the Starobinsky scalaron field. This provides a transition scale from a quadratic dominated gravity to a Einstein-Hilbert dominated gravity.
  • The model is well represented through a single mode scalar field characterized by a potential (11) (see Fig. 1) that can display the phenomenon of false vacuum decay, provided a condition is satisfied, already at classical level between the scalar Ricci curvature, that we assume constant, $\mathcal{R}$ satisfies the strong coupling condition $3M^{2}<\mathcal{R}$, being $M$ the mass term due to the quadratic term in the action (1).
  • The scalar field can be studied and quantized in the limit of a strong field dynamics. For $\Lambda=0$, this yields a mass gap and a spectrum represented in eq.(22). This is a discrete spectrum with an infinite tower of harmonic-oscillator like excitations. The spectrum depends on $G_{2}(x,x)$, a constant to be renormalized. The mass gap scale is proportional to the Ricci scalar $\mathcal{R}$.
  • We quantize the model using the technique of Dyson-Schwinger equations described in Appendix B as it is amenable to exact analytical solutions for the 1P- and 2P-correlation functions providing a Gaussian solution for the quantum field theory.
  • The quantized theory provides as a solution for the metric, the 1P-correlation function given in eq.(27), and representing the vacuum of the theory. A mass gap is displayed at the level of 2P-correlation function (31) that describes the propagating degrees of freedom.
  • The overall effect of a mass gap in the quadratic gravity makes it negligible at very low-energy where the EH term prevails.
Although we show the development of mass gap in the theory, if that were not the case, that is, we are not in the strong coupling regions and the conditions such as $\Lambda\gg M$ are not satisfied, there would be no way to observe such unusual features: simply because the patches of the Universe where their effects are sizable are not compatible with observers. It is only in the nearly homogeneous and isotropic Hubble patches that cosmic inflation may take place and experimental observations can eventually be done. This is one possible explicit quantum realization of the very popular Linde’s chaotic inflation Linde:1983gd and can be understood at the full quantum level of the classical results of quadratic gravity (for details, see Ref.Salvio:2019ewf ). This may naturally explain the near homogeneity and isotropy of our universe. We plan to study this using the mechanism presented in this paper and also extend this model adding matter content in future work.
Similarly, an interesting area of application of our formalism is for Einstein-Gauss-Bonnet cosmological models Nojiri:2005jg ; Cruz-DombrizS:2012 ; BenettiSCAL:2018 ; OdintsovOFF:2020 ; DeLaurentis:2015fea that appear very promising in the understanding of the experimental data Odintsov:2025kyw . We will address also this matter in future publications.
Acknowledgement
Authors thank Alexey Koshelev and Alberto Salvio for discussion.
Appendix A: Derivation of Dyson-Schwinger equations for a quartic scalar field
The technique we use to derive the set of Dyson-Schwinger equations in PDE shape has been devised by Bender, Milton and Savage Bender:1999ek for a PT-invariant quantum mechanical model. We extended this approach to quantum field theory Frasca:2015yva . For our convenience, we work with an Euclidean metric and write the partition function as
$Z[j]=\int[d\phi]e^{-\int d^{4}x[\frac{1}{2}(\partial\phi)^{2}-\frac{\lambda}{4}\phi^{4}+j\phi]}$ (33)
where $j$ is arbitrary source and $\lambda$ the theory self-coupling. Our aim is to evaluate the full set of nP-correlation functions
$G_{n}(x_{1},x_{2},\ldots,x_{n})=\left.\frac{\delta^{n}\ln Z}{\delta j(x_{1})\delta j(x_{2})\ldots\delta j(x_{n})}\right|_{j=0},$ (34)
so that we are able to get a complete solution for the quantum field theory and all the observables like decay rates and scattering cross sections will be computable through the LSZ theorem. It is straightforward to write down
$G_{1}(x_{1})$ $=$ $\left.\frac{1}{Z}\frac{\delta Z}{\delta j(x_{1})}\right|_{j=0},$
$G_{2}(x_{1},x_{2})$ $=$ $\left.\frac{\delta G_{1}^{(j)}(x_{1})}{\delta j(x_{2})}\right|_{j=0},$
$G_{3}(x_{1},x_{2},x_{3})$ $=$ $\left.\frac{\delta G_{2}^{(j)}(x_{1},x_{2})}{\delta j(x_{3})}\right|_{j=0},$ (35)
$\vdots$ ,
where
$G_{n}^{(j)}(x_{1},x_{2},\ldots,x_{n})=\frac{\delta^{n}\ln Z}{\delta j(x_{1})\delta j(x_{2})\ldots\delta j(x_{n})},$ (36)
that are the correlation functions dependent on the source $j$. The equation of motion takes the form
$\partial^{2}\phi+\lambda\phi^{3}=j.$ (37)
In order to get $G_{1}$, we average this equation with $Z$ obtaining
$\partial^{2}G_{1}^{(j)}(x)+\lambda Z^{-1}\langle\phi^{3}\rangle=j,$ (38)
but using
$ZG_{1}^{(j)}(x)=\langle\phi\rangle,$ (39)
we can iterate the derivative with respect to $j(x)$ obtaining
$ZG_{2}^{(j)}(x,x)+Z[G_{1}^{(j)}(x)]^{2}=\langle\phi^{2}\rangle,$
$3ZG_{1}^{(j)}(x)G_{2}^{(j)}(x,x)+ZG_{3}^{(j)}(x,x,x)+Z[G_{1}^{(j)}(x)]^{3}=\langle\phi^{3}\rangle.$ (40)
By direct substitution in eq.(38), we get the PDE for $G_{1}^{(j)}$
$\partial^{2}G_{1}^{(j)}(x)+\lambda\left\{3G_{2}^{(j)}(x,x)G_{1}^{(j)}(x)+G_{3}^{(j)}(x,x,x)+[G_{1}^{(j)}(x)]^{3}\right\}=j,$ (41)
and, by setting $j=0$, one has
$\partial^{2}G_{1}(x)+\lambda[3G_{1}(x)G_{2}(x,x)+G_{3}(x,x,x)+G_{1}^{3}(x)]=0.$ (42)
To get the equation for $G_{2}$. we need to derive eq.(41) with respect to $j(x_{2})$. This will yield
$\partial^{2}G_{2}^{(j)}(x,x_{2})+\lambda\left\{3G_{2}^{(j)}(x,x)G_{2}^{(j)}(x,x_{2})+3G_{2}^{(j)}(x,x)G_{3}^{(j)}(x,x,x_{2})+G_{4}^{(j)}(x,x,x,x_{2})+3[G_{1}^{(j)}(x)]^{2}G_{2}^{(j)}(x,x_{2})\right\}.$ (43)
By setting $j=0$, one gets the equation for $G_{2}$. The procedure can be iterated at any order to compute the higher-order correlation functions. This approach is successful when exact solution are known for $G_{1}$ and $G_{2}$. Only in this case, a Gaussian solution for the partition function can be obtained by selecting $G_{n}=0$ for $n>2$, provided at least two independent variables are identical.
Appendix B: Mass shift
We want to evaluate the contribution of the mass shift in eq.(30) by evaluating the self-consistency equation
$\mu_{R}^{2}=\mu_{0}^{2}-3\int\frac{d^{4}p}{(2\pi)^{4}}\frac{1}{p^{2}+2\mu_{R}^{2}}.$ (44)
Using dimensional regularization, we get
$\int\frac{d^{D}p}{(2\pi)^{4}}\frac{1}{p^{2}+2\mu_{R}^{2}}=\frac{\Gamma(-1+\epsilon)}{(4\pi)^{2-\epsilon}}(\mu_{R}^{2})^{1-\epsilon},$ (45)
where we set $\epsilon=(4-D)/2$. Thus, doing an expansion for $\epsilon\rightarrow 0$, we get
$\mu_{R}^{2}=\mu_{0}^{2}-3\left[-\frac{\mu_{R}^{2}}{16\pi^{2}\epsilon}+\frac{\mu_{R}^{2}}{16\pi^{2}}\left(-1+\gamma+\log(\mu_{R}^{2})+\log(4\pi)\right)\right]+O(\epsilon),$ (46)
where $\gamma$ is the Euler-Mascheroni constant. We take care of the divergent part by considering it a correction to the coupling constant $M$. Thus, we can safely take the limit $\epsilon\rightarrow 0$ and we are left with
$\mu_{R}^{2}\left[1+\frac{3}{16\pi^{2}}\left(-1+\gamma+\log(\mu_{R}^{2})+\log(4\pi)\right)\right]=\mu_{0}^{2}.$ (47)
This equation can be solved using the Lambert function $W$, writing
$\mu_{R}^{2}=\frac{4\pi^{2}\left(\sqrt{3}\sqrt{3M^{2}+8R}+3M\right)^{2}}{27W\left(\frac{16}{27}\pi^{3}\left(\sqrt{3}\sqrt{3M^{2}+8R}+3M\right)^{2}e^{-1+\gamma+\frac{16\pi^{2}}{3}}\right)}.$ (48)